OpenStudy (anonymous):

Identify the slope. y = 2.5x – 3.5 A. -5/2 B. 5/2 C. -3.5 D. 3.5

HanAkoSolo (jamierox4ev3r):

the slope is 2.5, if you take reference to the the equation; therefore the answer is b because 5/2=2.5

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

use y=mx+c where m is the slope

OpenStudy (anonymous):

isnt it y-mx+b?

OpenStudy (anonymous):

nope dear

HanAkoSolo (jamierox4ev3r):

y=mx+b lol